# CMB line-of-sight integrators for nearly-isotropic cosmological models

### v1

@@ I Introduction
A long-standing cosmological question concerns the global shape of the universe. According to the inflationary paradigm, our particle horizon is just a smooth patch of a larger manifold whose exact geometrical and topological properties are unknown. The extension of this patch beyond our particle horizon and the details of the underlying spacetime manifold are uncertain, and both have profound implications for inflation and the cosmological principle. At the same time, the existence of large-angle statistical anomalies in the cosmic microwave background (CMB) and persistent tensions between low- and high-redshift data [1, 2], suggest that new physics may be lurking just beyond the horizon. These considerations motivate the search for observational signatures that could reveal the global structure of spacetime while also testing the robustness of the concordance Friedmann-Lemaître-Robertson-Walker (FLRW) model. This program requires relaxing the foundational symmetry assumptions of our universe, namely, that it is spatially homogeneous and isotropic at large scales, and with topologically trivial boundary conditions. While past searches for nontrivial topologies resulted inconclusive [3, 4, 5], this is an active area of research that might be settled with existing and future cosmological data [6]. Another possibility is to keep the trivial topology, but change the geometrical setup. In this case the simplest option is to describe the universe using homogeneous and spatially anisotropic Bianchi models.
@@
In principle, any cosmological observable in a given nearly-isotropic model can be computed and compared against observations. In practice, however, this is a much more involved task than its isotropic analogous, and its success depends strongly on the type of observable chosen. Nonetheless, if spatial anisotropies are small, as observations suggest, one can show that the anisotropies will evolve as an independent set of modes on top of a maximally symmetric space—just as it happens with the usual scalar, vector and tensor modes of linear perturbation theory. This idea was originally introduced in [25, 26, 27], where homogeneous tensor perturbations of spherical FLRW universes were identified with the dynamics of Bianchi IX models. It was then generalized in [28] to include the proper definition of scalar, vector and tensor modes of nearly-isotropic models, and their identification with the dynamics of perturbative modes of FLRW spacetimes.
@@ II Homogeneous FLRW perturbations as Bianchi models
$\bm{g}^{\text{FLRW}}=-{\rm d}t^{2}+a^{2}(t)[\gamma_{ij}+2h_{ij}]{\rm d}x^{i}\otimes{\rm d}x^{j}\,,$ (1)
@@
$\bm{g}^{\text{Bianchi}}=-{\rm d}t^{2}+a^{2}(t)[\delta_{ij}+2\beta_{ij}]\bm{e}^{i}\otimes\bm{e}^{j}\,.$ (2)
@@
Naturally, this identification comes with important caveats. The first and most obvious is related to the choice of basis, since the tensor $h_{ij}$ is usually implemented in a coordinate basis $\{{\rm d}x^{i}\}$, whereas $\beta_{ij}$ is better described in terms of the so-called invariant basis $\{\bm{e}_{i}\}$, defined so as to respect the Bianchi symmetries. In particular, this distinction means that the homogeneous limit does not necessarily imply a homogeneous (i.e., position-independent) $h_{ij}$ [28]. Second, this identification can only hold for Bianchi models having maximally symmetric spaces as their isotropic limits. This includes models I and VII0 (having $\mathbb{E}^{3}$ as limit), models V and VIIh ($\mathbb{H}^{3}$), and model IX ($\mathbb{S}^{3}$). Finally, since $\beta_{ij}$ is traceless (because it is a volume-preserving deformation tensor), the trace of $h_{ij}$ does not play a role in the identification. [^1] The trace of $h_{ij}$ corresponds to a local curvature of the space in the homogeneous limit, and can be absorbed in a redefinition of the scale factor in Friedmann’s equation to account for this curvature — see [36].
@@
where ${\cal H}$ is the conformal Hubble factor and a prime means derivative with respect to $\eta=\int{\rm d}t/a$. This is dynamically equivalent to the equation for the symmetric and traceless shear tensor [^2] We caution the reader that we are reserving the term shear for the metric perturbation $\beta_{ij}$, while in most references this term is associated with the tensor ${\sigma_{ij}\equiv\beta^{\prime}_{ij}}$. $\beta_{ij}$ in a Bianchi-I spacetime:
@@
as one can easily check by computing the tracefree part of Einstein equations from (2) using $\bm{e}^{i}={\rm d}x^{i}$. An important difference in the two approaches is related to the choice of initial conditions. While in FLRW we are interested in the growing mode of $h_{ij}$, the growing mode of $\beta_{ij}$ is constant in this case, and thus a gauge artifact. Since decaying modes of $\beta_{ij}$ usually diverge towards the Big Bang, some care is needed to chose a non-trivial growing mode of the shear, as we will see.
In the general case, the identification (3) is done in terms of the modes of the perturbations that transform as scalars, vectors, or tensors under spatial rotations. For $h_{ij}$, it is usual to consider rotations around a fixed Fourier vector $\bm{\nu}$ when defining these modes, which leads to well-known scalar-vector-tensor (SVT) decomposition [^3] For closed FLRW universes, eigenfunctions are periodic, so that $\bm{\nu}=(\nu,\ell,M)$ is discrete, and the integral is replaced by a sum.
$h_{ij}=\sum_{m=-2}^{2}\int\frac{{\rm d}^{3}\bm{\nu}}{(2\pi)^{3}}h_{(m)}(\bm{\nu},\eta)Q^{(m)}_{ij}(\bm{\nu})\,,$ (6)
where $Q^{(m)}_{ij}$ are symmetric and traceless tensor harmonics of maximally symmetric 3-spaces [see Ref. [32] for their derivation] and $m=0$, ±$\pm 1$, and ±$\pm 2$ represent scalars, vectors or tensors perturbations, respectively.
Similarly, for $\beta_{ij}$, we can define the equivalent SVT modes, henceforth $svt$ modes, using rotations around any chosen vector of the invariant-basis defining a particular Bianchi model. Quite generally, it is convenient to consider $\bm{e}_{3}$ as a reference direction, and work with the so-called polarization basis $(\bm{e}^{(+)},\bm{e}^{(-)},\bm{e}_{3})$, with $\bm{e}^{(\pm)}=(\bm{e}_{1}\mp i\bm{e}_{2})/\sqrt{2}$. In this basis, $\beta_{ij}$ is decomposed as
$\beta_{ij}=\sum_{m=-2}^{2}\beta_{(m)}q^{(m)}_{ij}(\bm{e}_{3})\,,$ (7)
@@
$q^{(0)}_{ij}(\bm{e}_{3})$ $=(-e^{3}_{i}e^{3}_{j}+\delta_{ij}/3)\,,$ (8)
$q^{(\pm 1)}_{ij}(\bm{e}_{3})$ =±$=\pm e^{3}_{(i}e^{(\mp)}_{j)}\,,$ (9)
$q^{(\pm 2)}_{ij}(\bm{e}_{3})$ $=-\sqrt{3/2}\,e^{(\mp)}_{i}e^{(\mp)}_{j}\,.$ (10)
Finally, the proof of the equivalence (3) is done in two steps. First, from the linearized Einstein equations for the metrics (1) and (2), one looks for the wavenumber $\nu_{m}$ such that the SVT and $svt$ modes are dynamically equivalent. Second, for each such $\nu_{m}$, one verifies the equivalence of Eqs. (6) and (7), along with all their derivatives, at the origin of the coordinates. This leads to [29]
@@
and $\xi_{0}\equiv 1$. Note also the appearance of the flat-space wavenumber $k_{m}$, which is identical to $\nu_{m}$ when ${K}=0$, and otherwise given by [32]
@@
Type $K$ $\bm{\nu_{m}}$ $\bm{\zeta_{\ell}^{m}}$
I $=0$ 00 $\delta_{\ell}^{2}$
VII0 $=0$ $m/\ell_{s}$ (±$\left(\pm 1\right)^{\ell}$
IX $>0$$\pm 3/\ell_{c}$ $\delta_{\ell}^{2},\;|m|=2$ only
V $<0$ ${\rm i}/\ell_{c}$ (±$(\pm 1)^{m}\left(-{\rm i}\right)^{\ell-m}\sqrt{\frac{m(m+1)}{\ell(\ell+1)}}$
VIIh $<0$ $\frac{m}{\ell_{s}}+\frac{{\rm i}}{\ell_{c}}$ (±$(\pm 1)^{m}\!\prod_{p=m+1}^{\ell}(-{\rm i})\sqrt{\!\frac{(p-1)\sqrt{h}\pm{\rm i}m}{(p+1)\sqrt{h}\mp{\rm i}m}}$
@@ III CMB in the Line-of-sight approach
One of the main advantages of the identification between FLRW perturbations and nearly-isotropic models is the fact that observables of the latter can be computed from known expressions of the former, allowing for a unified description. In particular, the imprints of nearly-isotropic models on CMB anisotropies, which so far have been obtained using two separate codes, one for deterministic Bianchi anisotropies and another for stochastic FLRW perturbations [37, 15], can be unified into a single line-of-sight integrator [^4] See also [38] for a numerical implementation of temperature anisotropies in model VIIh considering only the Sachs-Wolfe effect.. This requires two main modifications in the structure of standard codes, one related to the Boltzmann hierarchy, and another to the domain of the radial Bessel functions entering the line-of-sight integrals.
@@ III.1 Boltzmann-Bianchi hierarchy
$\begin{split}{}_{s}\kappa^{m}_{\ell}&\rightarrow{}_{s}\kappa^{m}_{\ell}\frac{\zeta^{m}_{\ell}}{\zeta^{m}_{\ell-1}}\,,\\ {}_{s}\kappa^{m}_{\ell+1}&\rightarrow{}_{s}\kappa^{m}_{\ell+1}\frac{\zeta^{m}_{\ell}}{\zeta^{m}_{\ell+1}}\,,\end{split}$ (15)
@@
where $\tau$ is the optical depth, and the coefficients ${\cal G}^{(m)}_{\ell}$ and ${}^{X}{\cal C}^{(m)}_{\ell}$ ($X=T,E,B$) represent the gravitational and collisional terms, respectively. As we can see, the anisotropic dynamics have a direct impact on the free-streaming terms (inside square brackets) through the constants ${}_{s}\kappa^{m}_{\ell}$ and $\zeta^{m}_{\ell}$.
@@
$\begin{split}\beta^{\prime\prime}_{(\pm 2)}+2{\cal H}\beta^{\prime}_{(\pm 2)}&+(\nu^{2}_{\pm 2}-K)\beta_{(\pm 2)}=\\ &\frac{8a^{2}}{5}\left(\rho_{\gamma}\Theta^{(2)}_{2}+\rho_{\nu}{\cal N}^{(2)}_{2}\right),\end{split}$ (25)
@@
with $R=3{\rho}_{b}/4{\rho}_{\gamma}$. Finally, the shear equation for the vectors modes is
@@
$-\frac{8\sqrt{3}a^{2}}{5\sqrt{\nu^{2}_{\pm 1}-4K}}\left(\rho_{\nu}{\cal N}^{(\pm 1)}_{2}+\rho_{\gamma}\Theta^{(\pm 1)}_{2}\right)\,.$
@@ III.2 Radial functions
×$\qquad\qquad\qquad\times\frac{\zeta^{m}_{\ell}}{\zeta^{m}_{\ell^{\prime}}}{}_{0}\bar{\epsilon}_{\ell}^{(\ell^{\prime}m)}(\chi),$ (32)
@@ III.3 Initial conditions
An important difference between the usual SVT and the homogeneous $svt$ modes refers to their initial conditions. While the former are believed to result from some quantum random process during the early universe, the latter evolve from classical initial conditions. We now briefly discuss the initial conditions used to evolve $\beta_{(m)}$ and the baryon velocity $v_{b}$ in time in our numerical implementations.
@@
where ${\cal S}^{(m)}$ is a constant accounting for the spatial curvature—see [7, 29]. As remarked earlier, the solutions of the shear are classified as regular and irregular, according to their behavior at $\eta=0$. For ${\cal S}^{(m)}=0$, $\beta_{(m)}$ is either a pure constant, or divergent at $\eta=0$, so that these solutions are not cosmologically interesting. Since this is the case in model I, this model is not implemented in AniLoS or AniCLASS.
@@
They can be interpreted as superhorizon and non-stochastic gravitational waves which are frozen, but which become dynamical after entering the horizon. For example, in model VIIh and in a matter-dominated universe, one finds $\beta_{(\pm 2)}\propto j_{1}(\omega_{\pm}\eta)/(\omega_{\pm}\eta)$, where $\omega_{\pm}=\sqrt{1\pm i\ell_{s}/\ell_{c}}$. This solution can display regular solutions which are either constant or oscillating, depending on the values of the constants $\ell_{c}$ and $\ell_{s}$ [7].
@@
where we have used (14).
Regular vector modes ($|m|=1$) could also be observed in the presence of free-streaming neutrinos and isocurvature initial conditions. For example, an initial condition (after neutrino decoupling) where photons and neutrinos have non-zero and nearly opposite velocities leads to a constant vector mode [39, 46]. In the homogeneous limit, this corresponds to a vector part of the shear that does not decay, but remains constant on superhorizon scales at early times; in the codes, this conditions is labeled iso. Another possibility (labeled oct) is if neutrinos only have an initial octopole moment [47]. Through the Boltzmann equation, this octopole moment induces an anisotropic stress at early times, which can also lead to a regular vector mode of the shear. In AniLoS and AniCLASS, we implement the tensor regular solution and both isocurvature quadrupole and octopole initial conditions for vector modes.

### v2

@@ I Introduction
A long-standing cosmological question concerns the global shape of the universe. According to the inflationary paradigm, our particle horizon is just a smooth patch of a larger manifold whose exact geometrical and topological properties are unknown. The extension of this patch beyond our particle horizon and the details of the underlying spacetime manifold are uncertain, and both have profound implications for inflation and the cosmological principle. At the same time, the existence of large-angle statistical anomalies in the cosmic microwave background (CMB) and persistent tensions between low- and high-redshift data [10, 2], suggest that new physics may be lurking just beyond the horizon. These considerations motivate the search for observational signatures that could reveal the global structure of spacetime while also testing the robustness of the concordance Friedmann-Lemaître-Robertson-Walker (FLRW) model. This program requires relaxing some or all of the foundational symmetry assumptions of our universe, namely, that it is spatially homogeneous and isotropic at large scales (see [39, 26] for classical references and [22, 32] for recent accounts), and with topologically trivial boundary conditions [40]. While past searches for nontrivial topologies resulted inconclusive [6, 7, 45], this is an active area of research that might be settled with existing and future cosmological data [8]. Another possibility is to keep the trivial topology, but change the geometrical setup. In this case the simplest option is to describe the universe using homogeneous and spatially anisotropic Bianchi models.
@@
In principle, any cosmological observable in a given nearly-isotropic model can be computed and compared against observations. In practice, however, this is a much more involved task than its isotropic analogs, and its success depends strongly on the type of observable chosen. Nonetheless, if spatial anisotropies are small, as observations suggest, one can show that the anisotropies will evolve as an independent set of modes on top of a maximally symmetric space—just as it happens with the usual scalar, vector and tensor modes of linear perturbation theory. This idea was originally introduced in [24, 29, 37], where homogeneous tensor perturbations of spherical FLRW universes were identified with the dynamics of Bianchi IX models. It was then generalized in [60] to include the proper definition of scalar, vector and tensor modes of nearly-isotropic models, and their identification with the dynamics of perturbative modes of FLRW spacetimes.
@@ II Homogeneous FLRW perturbations as Bianchi models
$\boldsymbol{g}^{\text{FLRW}}=-{\rm d}t^{2}+a^{2}(t)[\gamma_{ij}+2h_{ij}]{\rm d}x^{i}\otimes{\rm d}x^{j}\,$, (1)
@@
$\boldsymbol{g}^{\text{Bianchi}}=-{\rm d}t^{2}+a^{2}(t)[\delta_{ij}+2\beta_{ij}]\boldsymbol{e}^{i}\otimes\boldsymbol{e}^{j}\,.$ (2)
@@
Naturally, this identification comes with important caveats. The first and most obvious is related to the choice of basis, since the tensor $h_{ij}$ is usually implemented in a coordinate basis $\{{\rm d}x^{i}\}$, whereas $\beta_{ij}$ is better described in terms of the so-called invariant basis $\{\boldsymbol{e}_{i}\}$, defined so as to respect the Bianchi symmetries. In particular, this distinction means that the homogeneous limit does not necessarily imply a homogeneous (i.e., position-independent) $h_{ij}$ [60]. Second, this identification holds for Bianchi models having maximally symmetric spaces as their isotropic limits. This includes models I and VII0 (having $\mathbb{E}^{3}$ as limit), models V and VIIh ($\mathbb{H}^{3}$), and model IX ($\mathbb{S}^{3}$). Models with anisotropic spatial geometries, also known as Thurston’s geometries [70], do not fall in this class, although they can also be constrained by CMB [53, 68] and distance measures [11]. Finally, since $\beta_{ij}$ is traceless (because it is a volume-preserving deformation tensor), the trace of $h_{ij}$ does not play a role in the identification. [^1] The trace of $h_{ij}$ corresponds to a local curvature of the space in the homogeneous limit, and can be absorbed in a redefinition of the scale factor in Friedmann’s equation to account for this curvature — see [50].
@@
[…]
@@
as one can easily check by computing the tracefree part of Einstein equations from (2) using $\boldsymbol{e}^{i}={\rm d}x^{i}$. An important difference in the two approaches is related to the choice of initial conditions. While in FLRW we are interested in the growing mode of $h_{ij}$, the growing mode of $\beta_{ij}$ is constant in this case, and thus a gauge artifact. Since decaying modes of $\beta_{ij}$ usually diverge towards the Big Bang, some care is needed to chose a non-trivial growing mode of the shear, as we will see.
In the general case, the identification (3) is done in terms of the modes of the perturbations that transform as scalars, vectors, or tensors under spatial rotations. For $h_{ij}$, it is usual to consider rotations around a fixed Fourier vector $\boldsymbol{\nu}$ when defining these modes, which leads to well-known scalar-vector-tensor (SVT) decomposition [^3] For closed FLRW universes, eigenfunctions are periodic, so that $\boldsymbol{\nu}=(\nu,\ell,M)$ is discrete, and the integral is replaced by a sum.
$h_{ij}=\sum_{m=-2}^{2}\int\frac{{\rm d}^{3}\boldsymbol{\nu}}{(2\pi)^{3}}h_{(m)}(\boldsymbol{\nu},\eta)Q^{(m)}_{ij}(\boldsymbol{\nu})\,$, (6)
where $Q^{(m)}_{ij}$ are symmetric and traceless tensor harmonics of maximally symmetric 3-spaces [see Ref. [59] for their derivation] and $m=0$, $\pm 1$, and $\pm 2$ represent scalars, vectors or tensors perturbations, respectively.
Similarly, for $\beta_{ij}$, we can define the equivalent SVT modes, henceforth svtsvt modes, using rotations around any chosen vector of the invariant-basis defining a particular Bianchi model. Quite generally, it is convenient to consider $\boldsymbol{e}_{3}$ as a reference direction, and work with the so-called polarization basis $(\boldsymbol{e}^{(+)},\boldsymbol{e}^{(-)},\boldsymbol{e}_{3})$, with $\boldsymbol{e}^{(\pm)}=(\boldsymbol{e}_{1}\mp i\boldsymbol{e}_{2})/\sqrt{2}$. In this basis, $\beta_{ij}$ is decomposed as
$\beta_{ij}=\sum_{m=-2}^{2}\beta_{(m)}q^{(m)}_{ij}(\boldsymbol{e}_{3})\,$, (7)
@@
$q^{(0)}_{ij}(\boldsymbol{e}_{3})$ $=(-e^{3}_{i}e^{3}_{j}+\delta_{ij}/3)\,$, (8)
$q^{(\pm 1)}_{ij}(\boldsymbol{e}_{3})$ $=\pm e^{3}_{(i}e^{(\mp)}_{j)}\,$, (9)
$q^{(\pm 2)}_{ij}(\boldsymbol{e}_{3})$ $=-\sqrt{3/2}\,e^{(\mp)}_{i}e^{(\mp)}_{j}\,$. (10)
Finally, the proof of the equivalence (3) is done in two steps. First, from the linearized Einstein equations for the metrics (1) and (2), one looks for the wavenumber $\nu_{m}$ such that the SVT and svtsvt modes are dynamically equivalent. Second, for each such $\nu_{m}$, one verifies the equivalence of Eqs. (6) and (7), along with all their derivatives, at the origin of the coordinates. This leads to [55]
@@
and $\xi_{0}\equiv 1$. Note also the appearance of the flat-space wavenumber $k_{m}$, which is identical to $\nu_{m}$ when K=0{K}=0, and otherwise given by [59]
@@
Type $K$ $\boldsymbol{\nu_{m}}$ $\boldsymbol{\zeta_{\ell}^{m}}$
I =0=0 0 $\delta_{\ell}^{2}$
VII0 =0=0 $m/\ell_{s}$ $\left(\pm 1\right)^{\ell}$
IX >0>0 $\pm 3/\ell_{c}$ $\delta_{\ell}^{2},\;|m|=2$ only
V <0<0 ${\rm i}/\ell_{c}$ $(\pm 1)^{m}\left(-{\rm i}\right)^{\ell-m}\sqrt{\frac{m(m+1)}{\ell(\ell+1)}}$
VIIh <0<0 $\frac{m}{\ell_{s}}+\frac{{\rm i}}{\ell_{c}}$ $(\pm 1)^{m}\!\prod_{p=m+1}^{\ell}(-{\rm i})\sqrt{\!\frac{(p-1)\sqrt{h}\pm{\rm i}m}{(p+1)\sqrt{h}\mp{\rm i}m}}$
@@ III CMB in the Line-of-sight approach
One of the main advantages of the identification between FLRW perturbations and nearly-isotropic models is the fact that observables of the latter can be computed from known expressions of the former, allowing for a unified description. In particular, the imprints of nearly-isotropic models on CMB anisotropies, which so far have been obtained using two separate codes, one for deterministic Bianchi anisotropies and another for stochastic FLRW perturbations [65, 66], can be unified into a single line-of-sight integrator. [^4] See also [48] for a numerical implementation of temperature anisotropies in model VIIh considering only the Sachs-Wolfe effect. This requires two main modifications in the structure of standard codes, one related to the Boltzmann hierarchy, and another to the domain of the radial Bessel functions entering the line-of-sight integrals.
@@ III.1 Boltzmann-Bianchi hierarchy
[…]
@@
where $\tau$ is the optical depth, and the coefficients ${\cal G}^{(m)}_{\ell}$ and $^{X}{\cal C}^{(m)}_{\ell}$ ($X={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Theta},E,B$) the gravitational and collisional terms, respectively. As we can see, the anisotropic dynamics have a direct impact on the free-streaming terms (inside square brackets) through the constants $_{s}\kappa^{m}_{\ell}$ and $\zeta^{m}_{\ell}$.
@@
[…]
@@
with $R=3{\rho}_{b}/4{\rho}_{\gamma}$. The shear equation for the vectors modes is
@@
$-\frac{8\sqrt{3}a^{2}}{5\sqrt{\nu^{2}_{\pm 1}-4K}}\left(\rho_{\gamma}\Theta^{(\pm 1)}_{2}+\rho_{\nu}{\cal N}^{(\pm 1)}_{2}\right)\,$.
@@ III.2 Radial functions
$\qquad\qquad\qquad\times\frac{\zeta^{m}_{\ell}}{\zeta^{m}_{\ell^{\prime}}}{}_{0}\bar{\epsilon}_{\ell}^{(\ell^{\prime}m)}(\chi),$ (32)
@@ III.3 Initial conditions
An important difference between the usual SVT and the homogeneous svtsvt modes refers to their initial conditions. While the former are believed to result from some quantum random process during the early universe, the latter evolve from classical initial conditions. We now briefly discuss the initial conditions used to evolve $\beta_{(m)}$ and the baryon velocity $v_{b}$ in time in our numerical implementations.
@@
where ${\cal S}^{(m)}$ is a constant accounting for the (anisotropic) spatial curvature—see [61, 55]. As remarked earlier, the solutions of the shear are classified as regular and irregular, according to their behavior at $\eta=0$. For ${\cal S}^{(m)}=0$, $\beta_{(m)}$ is either a pure constant, or divergent at $\eta=0$, so that these solutions are not cosmologically interesting. Since this is the case in model I, this model is not implemented in AniLoS or AniCLASS. The only exception is the tensor regular solution $\beta_{(\pm 2)}=\text{constant}$ of model V. However, this solutions is not independent, but rather given by the $\ell_{s}\rightarrow\infty$ limit of the (regular) tensor solution of model VIIh.
@@
They can be interpreted as superhorizon and non-stochastic gravitational waves which are frozen, but which become dynamical after entering the horizon. For example, in model VIIh and in a matter-dominated universe, one finds $\beta_{(\pm 2)}\propto j_{1}(\omega_{\pm}\eta)/(\omega_{\pm}\eta)$, where $\omega_{\pm}=2\ell_{s}^{-1}\sqrt{1\pm i\ell_{s}/\ell_{c}}$. This solution can display regular solutions which are either constant or oscillating, depending on the values of the constants $\ell_{c}$ and $\ell_{s}$ [61].
@@
where we have used (14). Note that the second linearly-independent solution of Eq. (37) can be irregular. In the example above, there is an irregular solution given by $\cos(\sqrt{-{\cal S}^{(\pm 2)}}\eta)/(\sqrt{-{\cal S}^{(\pm 2)}}\eta)$. We stress that irregular solutions are not interesting, and thus not implemented in the codes.
In the absence of sources and for adiabatic initial conditions, vectors modes ($|m|=1$), just like scalar modes, will only present irregular or constant solutions. However, regular and non-trivial solutions can be observed in the presence of free-streaming neutrinos and isocurvature initial conditions. For example, an initial condition (after neutrino decoupling) where photons and neutrinos have non-zero and nearly opposite velocities leads to a constant vector mode [44, 36]. In the homogeneous limit, this corresponds to a vector part of the shear that does not decay, but remains constant on superhorizon scales at early times; in the codes, this conditions is labeled iso. Another possibility (labeled oct) is if neutrinos only have an initial octopole moment [63]. Through the Boltzmann equation, this octopole moment induces an anisotropic stress at early times, which can also lead to a regular vector mode of the shear. In AniLoS and AniCLASS, we implement the tensor regular solution and both isocurvature quadrupole and octopole initial conditions for vector modes.